If jo's class has 24 kids in it and the ratio of girls to boys is 3:1, how many more girls are there than boys? 3:1, 6:2, 9:3, 12:4, 15:5, 18:6, 18+6=24, 18-6=12, so there are 12 more girls right? How do I get the answer in a simpler way?

Answer by rfer(16322)

3x+x=24
4x=24
x=24/4
x=6 boys
3x=18 girls
---------
18-6=12 more girls